Supplemental Operations Report March 16th

Supplemental Operations Report 16 Mar 2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Outpost water supply

Action taken for non-nominal systems: We shocked the system with bleach, drained as much as could be pumped out and refilled it halfway. It is not clear and there is additional material floating on the top. We are letting it settle overnight and will reevaluate tomorrow.

Generator: Still being manually run at night.

ScienceDome Dual Split: Off

Solar—Charge controllers off at night.

Solar— Nominal

Notes on power system: Nothing to report

Propane Reading, station tank – not noted %

Propane Reading, director tank— not noted %

Propane Reading, intern tank— not noted %

Propane Reading, generator— not noted %

Ethanol Free Gasoline – not known gallons.

Water (Outpost tank) – 50 gallons

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Still using Percy as well as Soji.

ATV’s Used: (Honda, 350.1, 350.2, 300): Honda and the 350 that runs (I always forget to note which one it is)

Reason for use: maintenance

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0.25 hours

Notes on ATVs: The headlights on the Honda to no work at all. I don’t know when this occurred but I only just noticed and confirmed it today.

HabCar used and why, where? Yes, to town for garbage and water runs

CrewCar used and why, where? No, it is parked in Grand Junction

Luna used and why, where? Yes, to town and for school field trips

Campus wide inspection, if action taken, what and why? Poop and MOOP patrols continue.

General notes and comments: Propane was delivered today, but I didn’t check the tank levels because of mud after a short storm this afternoon. The propane driver had run over the conduit in the narrow space between the campus and the outpost and I had put a barrier around it so I could move and rebury it, but he ran over the barrier and the conduit shattered. I will need to find a way to replace the pipe and protects the wires and move it in the next few days.

Summary of general operations: We are hosting two school groups this week. Each is here for two nights.

Summary of internet: I have not used the Starlink enough to evaluate it yet but my HughesNet account has no more data and three days until recharge, so Starlink will be my only internet until then and I will get a good look at it.

Summary of suits and radios: Nothing to report

Summary of Hab operations: There was some water leaking from the toilet on Monday, but it was limited and has not continued. It happened when the tank was full, so maybe that was the cause, although it appeared to be fresh water. No idea what happened.

Summary of Outpost operations: Intern trailer is being occupied by TMS ED James Burk.

Summary of GreenHab operations: The Supaero crews left a small growth chamber, which I cleaned and will install in the ScienceDome to grown microgreens for the remainder of the field season.

Summary of ScienceDome operations: ScienceDome has been used a lot this week for science experiences. We have been using the oven.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Repairs will be done on Friday.

Summary of health and safety issues: Nothing to report


Copyright © The Mars Society. All rights reserved. | Main Site